# pgfplots.statistics

### v1

@@ V Evaluation on the Coverage Control Problem
We evaluate our approach for asynchronous Perception-Action-Communication (PAC) on the coverage control problem [15] for a swarm of 32 robots in simulation. Coverage control requires the collaboration of a robot swarm to provide sensor coverage to monitor a phenomenon or features of interest in an environment. An importance density function (IDF) [5] is used to model the probability distribution of features of interest. The problem is widely studied in robotics and has applications in various domains, including mobile networking [16], surveillance [17], and target tracking [18]. We consider the decentralized setup where the robots have limited sensing and communication capabilities. Furthermore, the environment is not known a priori, and the robots use their sensor to make localized observations of the environment.
@@
We model a 1024 m×$\times$1024 m rectangular environment and robots that make 64 m×$\times$64 m localized sensor observations. Each robot maintains a local map of size 256 m×$\times$256 m with cumulatively added observations and an obstacle map of the same size for the boundaries of the environment. Based on our contemporary work on coverage control, these maps are processed by a CNN with three layers of 32 latent size each. Additionally, relative positions of neighboring robots, obtained either through sensing or communication, within a communication radius of 128 m is mapped onto a two-channel ($x$ and $y$) normalized heatmap. These four maps are concatenated to form the input to a CNN, which is composed of three layers of 32 latent size and generates a 32 dimensional feature vector for each robot. The sensing and processing of maps by CNN constitute the perception module of the PAC framework. The output of the CNN is augmented with the normalized position of the robot to form a 34-dimensional feature vector, which is used as the input to the GNN with two layers of 256 latent size and $K=3$ hops of communication. The output of the GNN is processed by a multi-layer perceptron (MLP), i.e., an action NN, with one hidden layer and 32 latent size to compute the control velocities for the robot. These CNN and GNN architectures are part of an ongoing work on coverage control [^1] Appropriate references will be added in the final publication..
The entire model is trained using imitation learning with the ground truth generated using a centralized clairvoyant Lloyd’s algorithm that has complete knowledge of the IDF and the positions of the robots. We compare our approach with a decentralized and a centralized Lloyd’s algorithm [19]. The decentralized algorithm is only aware of its maps and the relative positions of neighboring robots, whereas the centralized algorithm has the combined knowledge of all the robots and their global positions. We executed these algorithms in our asynchronous PAC framework for four different operating frequencies from (1.25 Hz to 5 Hz) of the perception module. Our approach performs significantly better than the decentralized and centralized algorithms, see Figure 4. We also evaluated our approach with noisy position information, see Figure 5. These results show the efficacy of asynchronous PAC systems for decentralized control of robot swarms.
VI Conclusion
@@
We presented a framework for asynchronous Perception-Action-Communication with Graph Neural Networks (GNNs). Perception comprises application-specific tasks such as sensing and mapping, while the action module is responsible for executing controls on the robot. The GNN bridges these two modules and provides learned messages for collaboration with other agents. The decentralized nature of the GNN enables scaling to large robot swarms. The asynchronous capability of our system allows the execution of GNN-based message aggregation and inferencing at a frequency in between that of perception and action modules. We demonstrated the effectiveness of using learnable PAC policies in a decentralized manner for the coverage control problem. The framework will allow evaluation and deployment of asynchronous PAC systems with GNNs with large robot swarms in real-world applications.
@@
Future work entails validating the system on real robots—an essential yet challenging task due to difficulties operating a large number of robots. We also plan to evaluate our framework on other applications, such as flocking and target tracking, and analyze the compatibility of our framework with other GNN architectures.

### v2

@@ VII Real-World Coverage Experiments
Coverage control requires the collaboration of a robot swarm to provide sensor coverage to monitor a phenomenon or features of interest in an environment. An importance density function (IDF) [5] is used to model the probability distribution of features of interest. The problem is widely studied in robotics and has applications in various domains, including mobile networking [22], surveillance [23], and target tracking [24]. We consider the asynchronous setup where the robots have limited sensing capabilities. Furthermore, the environment is not known a priori, and the robots use their sensor to make localized observations of the environment.
@@
We model a $\times1024 m$ rectangular environment and robots that make $\times64 m$ localized sensor observations. Each robot maintains a local map of size $\times256 m$ with cumulatively added observations and an obstacle map of the same size for the boundaries of the environment. Based on our contemporary work on coverage control, these maps are processed by a CNN with three layers of 32 latent sizes each. Additionally, relative positions of neighboring robots, obtained either through sensing or communication, within a communication radius of 128 m is mapped onto a two-channel ($x$ and $y$) normalized heatmap. These four maps are concatenated to form the input to a CNN, which is composed of three layers of 32 latent sizes and generates a 32-dimensional feature vector for each robot. The sensing and processing of maps by CNN constitute the perception module of the PAC framework. The output of the CNN is augmented with the normalized position of the robot to form a 34-dimensional feature vector, which is used as the input to the GNN with two layers of 256 latent size and $K=3$ hops of communication. The output of the GNN is processed by a multi-layer perceptron (MLP), i.e., an action NN, with one hidden layer and 32 latent sizes to compute the control velocities for the robot. These CNN and GNN architectures originate from [27].
The entire model is trained using imitation learning with the ground truth generated using a centralized clairvoyant Lloyd’s algorithm that has complete knowledge of the IDF and the positions of the robots. We evaluate our approach for asynchronous Perception-Action-Communication (PAC) on the coverage control problem [28] for a swarm of 10 robots in the real world. Our experiments are conducted outdoors with a team of 10 autonomous aerial vehicles (robot). The terrain is grassy, flat and contains no obstacles within the defined mission boundaries. Ten elevated launch pads are distributed uniformly throughout the environment. Each robot is assigned to a launch pad which is maintained throughout the experiments. We did this to this ensure consistency between experiments, as each robot has slight hardware peculiarities that make it unique. During our experiments, there were some instances of robot failure. In this case, we substituted an identical robot from our reserve to complete the remaining experiments.
Evaluation in the real-world was carried out in two experiments (E1) Transferability and (E2) Repeatability. (E1) tests the system’s performance under three different randomly generated IDFs; each with 10 Gaussian features. The purpose of (E2) is to assess the system’s behavior when subjected to the high variance of the real world. Localization, relying primarily on GPS, is especially susceptible to noise. The localization error can vary between robots as a result of variances in hardware and their location. For example, some robots have stronger GPS reception than others. Often, robots must operate with horizontal positional error in the range of $1-$5\text{\,}\mathrm{m}$\$. For (E3), during each experiment, we intentionally perturb the system by taking manual control of one robot. The robot is removed from the mission area, before it is manually returned. We only initiate perturbation once the system appears to have reached steady state. We always perturb the same drone and return it to approximately the same location near the mission boundary.
The original LPAC model was trained on an environment size of $1024\times 1024$ with $N$ robots and $F$ Gaussian features where $N=F=32$. There, the density of robots was $3.1e-5\;\text{robots}/\text{cell}$. For our experiments we selected a similar density of $3.8e-4\;\text{robots}/\text{cell}$ by reducing the environment size to $512\times 512$ and using $N=F=10$. We maintain this environment size and density for all experiments. Deploying the system on an environment of size 512×512 $\times$512\text{\,}{\mathrm{m}}^{2}$\$ is impractical for iterative development and experimentation. Robot positions are scaled by a factor of $16\;${\mathrm{m}}^{-1}$\$ to obtain a practical mission area of 32×32 $\times$32\text{\,}{\mathrm{m}}^{2}$\$. The scale factor is applied between the LPAC model and the offboard controller. This enables us to operate LPAC closer to the environment size used during training while deploying in a confined area.
The maximum output speed of the model is $5\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$. To deploy in a smaller environment, we reduce the velocity using a scale factor of $0.25$, resulting in a maximum speed of $1.25\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.
VII-A Transferability
A key boon of LPAC is its transferability to new environments. We verify that this property still holds in the real world by deploying our system in three unknown environments outside of the training set. The start locations for each robot are the same across all three environments.
Snapshots from these experiments are shown in Figure 8. In all three environments, the LPAC policy achieves coverage of the importance field. The global map is overlayed with low opacity over each image of the map. This is purely for visualization and to demonstrate the efficacy of LPAC. The robots do not have access to any information outside of their local observations and embeddings from neighbors.
Env 1 ($N=F=10$)
Env 2 ($N=F=10$)
Env 3 ($N=F=10$)
$0\text{\,}\mathrm{s}$ $15\text{\,}\mathrm{s}$ $30\text{\,}\mathrm{s}$ $60\text{\,}\mathrm{s}$
Figure 8: Performance of the LPAC system in the real world across three different environments in a 32×32 $\times$32\text{\,}{\mathrm{m}}^{2}$\$ area. Each environment contains $N=F=10$ robots and Gaussian features. Convergence typically occurs within $60\text{\,}\mathrm{s}$ and is dependent on environment and initial positions of the robots. The $\times$ markers denote the positions of the robots in the environment, while the IDF is illustrated in red. Each robot has a 4×4 $\times$4\text{\,}{\mathrm{m}}^{2}$\$ sensor area with which it explores the environment. The environment outside this sensor area is initially unknown to the agents, colored white here. We include a low opacity grayscale rendering of the IDF in unexplored regions to better convey coverage.
VII-B Repeatability
Under ideal, noiseless conditions, LPAC is a deterministic control policy. The real-world contains noise. We study its effect on LPAC by running the same experiment three times and observing the behavior of the system. Each experiment is run with robots starting from the same launch positions (a placement error within $10\text{\,}\mathrm{cm}$). The cost over time is plotted in Figure 9. The three cost traces capture the normalized coverage cost of each experiment within a window of $40\text{\,}\mathrm{s}$ from takeoff. This allows the system to converge to a stable coverage cost. The final cost achieved in all three experiments is similar despite differences in convergence rate. Despite physically placing the robots on the same launch pads, the localization error is significant. This error is magnified when the robots are near the ground and GNSS suffers from signal blockage and multipath interference. We capture the differences in initial and final positions between each of the three trials in Figure 10.
Figure 9: Evaluating repeatability. The normalized coverage cost of three trials captured over $40$ seconds. The beginning of each experiment occurs when the takeoff command is received. The trajectory of the cost varies between experiments but all ultimately reach a similar coverage cost. Factors such as noisy localization and wind play a significant role in how the quickly the robots converge to a solution in each trial.
(a) $t=$0\text{\,}\mathrm{s}$\$
(b) $t=$40\text{\,}\mathrm{s}$\$
Figure 10: Despite physically launching from the same site in each trial, the localization error in the robots is significant (left). The final reported local positions after $40\text{\,}\mathrm{s}$ (right) demonstrates the spread of positions once in the final timestep. A global view of the IDF is shown.
VIII Conclusion
@@
We presented a framework for asynchronous Perception-Action-Communication with Graph Neural Networks (GNNs). Perception comprises application-specific tasks such as sensing and mapping, while the action module is responsible for executing controls on the robot. The GNN bridges these two modules and provides learned messages for collaboration with other agents. The decentralized nature of the GNN enables scaling to large robot swarms. The asynchronous capability of our system allows the execution of GNN-based message aggregation and inferencing at a frequency in between that of perception and action modules. We demonstrated the effectiveness of using learnable PAC policies in a decentralized manner for real-world incarnations of the coverage control problem. The framework is scalable and allows the evaluation and deployment of asynchronous PAC systems with GNNs with large robot swarms.
@@
Future work entails validating the system with even larger scale teams and decentralized communications. We also plan to evaluate our framework on other applications, such as flocking and target tracking, and analyze the compatibility of our framework with other GNN architectures.